The Average weight of 25 chocolate bars selected from a normally distributed population is 200 grams with a standard deviation of 10 grams. Find the point estimate and the confidence interval using 95%confidence level.


Steps: Solution:

3.Collect and present sample evidence.

a.Collect the sample information The sample information consists of

_____(6),_______(7)and______(8).

b.Find the point estimate. The point estimate for the population

mean is ____________(9)


1
Expert's answer
2022-04-26T10:27:04-0400

The sample information consists of mean 200 g, standard deviation 10g and number of sample 25 chocolate.

The point estimate for the population


mean is 200 g

Z(95%) =1.96 from z value table

"CI=\mu \plusmn z\frac{\sigma}{\sqrt{n}}=200 \plusmn 1.96\frac{10}{\sqrt{25}}=200 \plusmn 3.92"